An investment of $93,000 was made by a business club. the investment was split into three parts and lasted for one year. The first part of the investment earned 8% interest, the second earned 6% and the third 9%. total interest from the investments was $7110. The interest from the first investment was 3 times the interest from the second. Find the amounts of the three parts of th investment??
Answer #1
second part --- x
first part ---- 3x
third part --- (93000 - x - 3x) = 93000 - 4x

.08(3x) + .06x + .09(93000-4x) = 7110
.24x + .06x + 8370 - .36x = 7110
-.06x = -1260
x = -1260/-.06 = 21000

second investment = x = $21000
first investment = 3x = $63000
third investment = 93000-4x = $9000


check:
63000(.08) + 21000(.06) + 9000(.09) = 7110
answered by: not telling